# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 4.  Balanced intervals and certain bijective maps
We start by defining balancedness of intervals in a slightly more general setting because this will help make the arguments clearer. Note that, as before, all intervals are understood modulo $1$. To make this more rigorous, we speak of the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, which can be thought of as the interval $[0,1)$, where we compute modulo $1$.
@@
Let $A=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}$ be a set of $m$ distinct points on the torus $\mathbb{T}$ and let $\delta\in(0,1)$. We say that the intervals of length $\delta$ are balanced with respect to $A$ if there exists an integer $c=c(A,\delta)$ such that for all $\xi\in\mathbb{T}$ we have
$([\xi,\xi+\delta)\cap A)\in\{c,c+1\}.$
The goal of this section is to find a way to determine this balancedness property without actually counting the points in each interval (this morally corresponds to the idea in [7, Lemma 1]). The first step is to focus on the intervals $\xi_{\ell},\xi_{\ell}+\delta)$ for $0\leq\ell\leq m-1$ and to find the closest points $\xi_{j}\in A$ to the left and to the right of $\xi_{\ell}+\delta$. We define the corresponding functions in terms of the indices of the points.
@@
Let $A=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points and let $\delta\in(0,1)$. Then we define the following two maps on the set $\{0,1,\ldots,m-1\}$: $f_{\textnormal{left}}$ maps $\ell$ to the index of the closest point in $A$ that lies to the left of $\xi_{\ell}+\delta$, and $f_{\textnormal{right}}$ maps $\ell$ to the index of the closest point in $A$ that lies to the right of $\xi_{\ell}+\delta$. In other words,
$f_{\textnormal{left}}(\ell)=j\quad$ $:\iff\quad\xi_{\ell}+\delta-\xi_{j}=\min_{0\leq i\leq m-1}\xi_{\ell}+\delta-\xi_{i};$
@@
Now in the case that the interval length $\delta$ does not match the distance between any two points in $A$, we have the following main lemma.
@@
Let $A=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points. Moreover, let $\delta\in(0,1)$ be such that $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$. Then the following statements are equivalent:
@@
    The intervals of length $\delta$ are balanced with respect to $A$.
@@
The assumption in Lemma 4.3 that $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$ is necessary. For example, consider the three points $\xi_{0}=0$, $\xi_{1}=1/4$, and $\xi_{2}=1/2$, and set $\delta=1/2$. Then it is easy to check that the half open intervals of length $\delta$ either contain $1$ or $2$ points, and thus are balanced. However, $f_{\textnormal{left}}(0)=2=f_{\textnormal{left}}(1)$ and $f_{\textnormal{right}}(2)=0=f_{\textnormal{right}}(1)$. One can also check that changing the definitions of $f_{\textnormal{left}},f_{\textnormal{left}}$ to contain the restriction “strictly to the left/right” does not resolve the issue.
@@
$A=\{0,\alpha,\ldots,(m-1)\alpha\}\subseteq\mathbb{T}$
@@
Let $A=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points and let $\delta\in(0,1)$. Then the intervals of length $\delta$ are balanced with respect to $A$ if and only if the intervals of length $1-\delta$ are balanced. Moreover, if $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$, then for all $\varepsilon$ with $\lvert\varepsilon\rvert$ sufficiently small, the intervals of length $\delta$ are balanced if and only if the intervals of length $1-\delta+\varepsilon$ are balanced.
@@
The first statement is clear because the half open intervals of length $1-\delta$ correspond to the complements of the intervals of length $\delta$. If $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$, then it is clear that for $\lvert\varepsilon\rvert<\min_{i,j}\lvert\delta-(\xi_{i}-\xi_{j})\rvert$ we can modify the interval length $\delta$ to $\delta+\varepsilon$ without changing the occurring values of $([\xi,\xi+\delta)\cap A)$. ∎
@@ 5.  Properties of convergents and Ostrowski representations
The strategy for proving our main result (Theorem 2.1) in the next section will be to apply Lemma 4.3 to the points $\xi_{\ell}=\left\{\ell\alpha\right\}$ with $0\leq\ell\leq m-1$ and $\delta=\left\{n\alpha\right\}$. Therefore, the right endpoints of the intervals $\xi_{\ell},\xi_{\ell}+\delta)$ will be of the shape $\left\{\ell\alpha\right\}+\left\{n\alpha\right\}=\left\{(\ell+n)\alpha\right\}$, where $\ell+n>m-1$. In order to determine $f_{\textnormal{left}}(\ell),f_{\textnormal{right}}(\ell)$ we will need to figure out which point $\left\{j\alpha\right\}$ with $0\leq j\leq m-1$ lies closest to the left and which one closest to the right of $\left\{(\ell+n)\alpha\right\}$. We will do this using the Ostrowski representation mentioned in Section 2.
We now recall several properties of Ostrowski representations which will be useful.
All lemmas below are probably well known to experts (except perhaps for Lemma 5.17, but this one is not very hard to see either). Rather unfortunately though, there does not seem to exist a suitable reference book. In order to provide proofs for the lemmas, we recall some basic properties of convergents first; see, e.g., [6] for a reference. Throughout the rest of this paper, $\alpha\in(0,1)$ is a fixed irrational, and all Ostrowski representations are with respect to $\alpha$.
@@ 5.1.  Basic properties of convergents
As before, let $p_{k}/q_{k}$ denote the convergents to $\alpha\in(0,1)$. Then the numerators and denominators follow the recursions
@@
Then we have (see, e.g., [6, p. 9])
(5.3) $\delta_{k}=(-1)^{k}\lVert q_{k}\alpha\rVert\quad\text{for }k\geq 1\quad\text{(or }k\geq 0\text{, if }\alpha<1/2\text{)},$
@@ 5.2.  Basic properties of Ostrowski representations
Moreover, note that if $a_{1}=1$ (which is equivalent to $\left\{\alpha\right\}>1/2$, and to $q_{1}=q_{0}=1$), then $b_{0}=0$, and the lowest possible nonzero term in the representation is $b_{1}q_{1}$. This, and the fact that $\lVert q_{0}\alpha\rVert<1/2$ only holds if $\alpha<1/2$, is why we will often have to distinguish between the two cases $\alpha<1/2$ and $\alpha>1/2$.
@@
The next lemma describes the fact that if “$n$ contains no small $q_{k}$’s”, then $n\alpha$ is close to 0 modulo $1$. In particular, this means that digits with large index only have little impact on the position of $\left\{n\alpha\right\}$. In this paper, we are mostly interested in precise bounds for the size of $\lVert n\alpha\rVert$ if $n$ only has large digits.
@@
    $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$, provided that $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$).
Some other, more precise estimates, are the following:
@@
    $\lVert n\alpha\rVert>\lVert q_{L+1}\alpha\rVert+\lVert q_{N+2}\rVert$, if $n\leq q_{N}$;
@@
    $\lVert n\alpha\rVert>\lVert q_{L}\alpha\rVert+\lVert q_{L+1}\alpha\rVert$ if $b_{L}(n)\geq 2$ and $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$);
@@
    $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert-\lVert q_{N+2}\rVert$, provided that $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$), and $n\leq q_{N}$;
@@
    $\lVert n\alpha\rVert<\lVert(q_{L-1}+q_{L})\alpha\rVert$ if $b_{L}(n)\leq a_{L+1}-1$ and $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$);
@@
    $\lVert n\alpha\rVert<\lVert n^{\prime}\alpha\rVert$ if $k_{0}(n^{\prime})=k_{0}(n)=L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$) and $b_{L}(n)<b_{L}(n^{\prime})$;
@@
    $\lVert n\alpha\rVert<\lVert q_{L}\alpha\rVert$ if and only if the Ostrowski representation of $n$ is of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+1+2t}>0$.
@@
All statements follow pretty straightforwardly from (5.3) and (5.7). See also, for example, [6, p. 24–25, Lemma 1 and Theorem 1]. ∎
@@
Let $n\geq 1$ and assume that $k_{0}(n)\geq 2$ (or $k_{0}(n)\geq 1$ if $\alpha<1/2$). Then we have
@@
Proof.
This follows from the previous statements. See, for example, [3, Lemma 4.1] for a reference. ∎
The above lemma cannot be extended to $k_{0}(n)\geq 1$ (or $k_{0}(n)\geq 0$ if $\alpha<1/2$), so when the smallest allowed digit in the representation of $n$ shows up, then $\left\{n\alpha\right\}$ might lie on “the wrong side of 1/21/2”. For example, if $\alpha<1/2$ and $k_{0}(n)=0$, we might have $\left\{n\alpha\right\}>1/2$, even though from the parity of $k_{0}(n)$ we would expect $\left\{n\alpha\right\}<1/2$.
The next lemma will be useful when dealing with such exceptional cases.
@@
Analogously, if $\alpha>1/2$ and $k_{0}(n)=1$, then
$\left\{n\alpha\right\}<1/2\quad\implies\quad\left\{(n+1)\alpha\right\}<\left\{n\alpha\right\}.$
Note that in both cases the implication is that $\lVert(n+1)\alpha\rVert<\lVert n\alpha\rVert$.
@@
Assume that $\alpha<1/2$ and that $n=\sum_{k=0}^{N}b_{k}q_{k}$ with $b_{0}>0$. Then we get from (5.3) and (5.7) that
@@
Thus, $\left\{n\alpha\right\}+\left\{\alpha\right\}<1$ and so $\left\{n\alpha\right\}<\left\{(n+1)\alpha\right\}$, as desired.
The proof for the case $\alpha>1/2$ is analogous. ∎
@@ 5.3.  One-sided approximations
The overload of notation is justified by the fact that we will go on to focus on one of $\langle\cdot\rangle_{\textnormal{left}}$, $\langle\cdot\rangle_{\textnormal{right}}$ at a time; see also the lemma below. Moreover, in Section 6 we will consider the functions $f_{\textnormal{left}},f_{\textnormal{right}},x_{\textnormal{left}},x_{\textnormal{right}}$, and the definition above will be compatible with those functions. Note that if $\{\langle\cdot\rangle,\langle\cdot\rangle^{\prime}\}=\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$, we have
@@
$\langle\xi_{1}\rangle+\langle\xi_{2}\rangle<1\quad\iff\quad\langle\xi_{1}+\xi_{2}\rangle=\langle\xi_{1}\rangle+\langle\xi_{2}\rangle;$
$\langle\xi_{1}\rangle\geq\langle\xi_{2}\rangle\quad\iff\quad\langle\xi_{1}-\xi_{2}\rangle=\langle\xi_{1}\rangle-\langle\xi_{2}\rangle.$
@@
$\langle\xi\rangle<\langle\xi+\xi_{1}\rangle<\langle\xi+\xi_{2}\rangle\quad\implies\quad\langle\xi_{1}\rangle<\langle\xi_{2}\rangle.$
We start with a simple lemma for the situation where $\langle n\alpha\rangle>1/2$.
@@
Let $\langle\cdot\rangle\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$. If $\langle n\alpha\rangle>1/2$, then either $\langle(n-1)\alpha\rangle<\langle n\alpha\rangle$ or $\langle(n+1)\alpha\rangle<\langle n\alpha\rangle$.
@@
Since either $\langle\alpha\rangle<1/2$ or $\langle-\alpha\rangle<1/2$, we have that $\langle\pm\alpha\rangle<1/2<\langle n\alpha\rangle$ for one choice of $\pm$. Then $\langle(n\mp 1)\alpha\rangle=\langle n\alpha\rangle-\langle\pm\alpha\rangle<\langle n\alpha\rangle$. ∎
@@
(5.8) $\langle\cdot\rangle:=\begin{cases}\langle\cdot\rangle_{\textnormal{left}},&\text{if }\lVert n\alpha\rVert=\left\{n\alpha\right\};\\ \langle\cdot\rangle_{\textnormal{right}},&\text{if }\lVert n\alpha\rVert=\left\{-n\alpha\right\}.\end{cases}$
In other words, we choose $\langle\cdot\rangle$ such that $\langle n\alpha\rangle=\lVert n\alpha\rVert<1/2$.
We rephrase Lemma 5.2 in terms of $\langle\cdot\rangle$:
@@
Let $n,n^{\prime}\geq 1$ with $k_{0}(n),k_{0}(n^{\prime})\geq 2$ (or $\geq 1$ if $\alpha<1/2$). Set $\langle\cdot\rangle$ as in (5.8). Then
$\langle n^{\prime}\alpha\rangle<1/2\quad\iff\quad k_{0}(n)\equiv k_{0}(n^{\prime})\pmod{2}$.
The next two lemmas describe how we can compare $\langle n\alpha\rangle$ and $\langle n^{\prime}\alpha\rangle$ by looking at the lowest nonzero digits in $n,n^{\prime}$.
@@
Let $n\geq 1$ and set $\langle\cdot\rangle$ as in (5.8). Then for all $n^{\prime}\geq 1$ we have
$k_{0}(n^{\prime})<k_{0}(n)\quad\implies\quad\langle n\alpha\rangle<\langle n^{\prime}\alpha\rangle.$
@@
Assume first that $k_{0}(n)\geq 3$ (or $k_{0}(n)\geq 2$ if $\alpha<1/2$). Then if $k_{0}(n^{\prime})=k_{0}(n)-1$, we have from Lemma 5.5 that $\langle n^{\prime}\alpha\rangle>1/2>\langle n\alpha\rangle$. If $k_{0}(n^{\prime})\leq k_{0}(n)-2$, then we obtain $\langle n^{\prime}\alpha\rangle>\langle n\alpha\rangle$ from Lemma 5.1(1, 2).
If $k_{0}(n)=0$ (or $k_{0}(n)=1$ if $\alpha>1/2$), the implication is trivial.
We are left with the cases $k_{0}(n)=1$ if $\alpha<1/2$, and $k_{0}(n)=2$ if $\alpha>1/2$. We focus on $k_{0}(n)=1$ and $\alpha<1/2$ (the other case is analogous). Lemma 5.2 implies that $\langle\cdot\rangle=\langle\cdot\rangle_{\textnormal{right}}$. Assume $k_{0}(n^{\prime})=0$. If $\left\{n^{\prime}\alpha\right\}<1/2$, then $\langle n^{\prime}\alpha\rangle_{\textnormal{right}}>1/2>\langle n\alpha\rangle_{\textnormal{right}}$, as desired. If $\left\{n^{\prime}\alpha\right\}>1/2$, then as in the proof of Lemma 5.3 we see that $\left\{n^{\prime}\alpha\right\}<1-\left\{\alpha\right\}$, and so $\langle n^{\prime}\alpha\rangle_{\textnormal{right}}>\lVert\alpha\rVert=\lVert q_{0}\alpha\rVert>\lVert n\alpha\rVert=\langle n\alpha\rangle_{\textnormal{right}}$, where we used Lemma 5.1(2). ∎
@@
Let $n,n^{\prime}\geq 1$ with $k_{0}(n)=k_{0}(n^{\prime})=L$. Assume $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$). Set $\langle\cdot\rangle$ as in (5.8). Then
$b_{L}(n)<b_{L}(n^{\prime})\quad\implies\quad\langle n\alpha\rangle<\langle n^{\prime}\alpha\rangle.$

### v2

@@ 4.  Balanced intervals and certain bijective maps
We start by defining balancedness of intervals in a slightly more general setting because this makes the arguments clearer. Note that, as before, all intervals are understood modulo $1$. To make this more rigorous, we speak of the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, which can be thought of as the interval $[0,1)$, where we compute modulo $1$.
@@
Let $B=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}$ be a set of $m$ distinct points on the torus $\mathbb{T}$ and let $\delta\in(0,1)$. We say that the intervals of length $\delta$ are balanced with respect to $B$ if there exists an integer $c=c(B,\delta)$ such that for all $\xi\in\mathbb{T}$ we have
$([\xi,\xi+\delta)\cap B)\in\{c,c+1\}.$
The goal of this section is to find a way to determine this balance property without actually counting the points in each interval. This morally corresponds to the idea in [8, Lemma 1]. The first step is to focus on the intervals $\xi_{\ell},\xi_{\ell}+\delta)$ for $0\leq\ell\leq m-1$ and to find the closest points $\xi_{j}\in B$ to the left and to the right of $\xi_{\ell}+\delta$. We define the corresponding functions in terms of the indices of the points.
@@
Let $B=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points and let $\delta\in(0,1)$. Then we define the following two maps on the set $\{0,1,\ldots,m-1\}$: $f_{\textnormal{left}}$ maps $\ell$ to the index of the closest point in $B$ that lies to the left of $\xi_{\ell}+\delta$, and $f_{\textnormal{right}}$ maps $\ell$ to the index of the closest point in $B$ that lies to the right of $\xi_{\ell}+\delta$. In other words,
$f_{\textnormal{left}}(\ell)=j\quad$ $:\iff\quad\xi_{\ell}+\delta-\xi_{j}=\min_{0\leq i\leq m-1}(\xi_{\ell}+\delta)-\xi_{i},$
@@
Now in the case that the interval length $\delta$ does not match the distance between any two points in $B$, we have the following main lemma.
@@
Let $B=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points. Moreover, let $\delta\in(0,1)$ be such that $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$. Then the following statements are equivalent:
@@
    The intervals of length $\delta$ are balanced with respect to $B$.
@@
The assumption in Lemma 4.3 that $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$ is necessary. For example, consider the three points $\xi_{0}=0$, $\xi_{1}=1/4$, and $\xi_{2}=1/2$, and set $\delta=1/2$. Then it is easy to check that the half open intervals of length $\delta$ contain either $1$ or $2$ points, and thus are balanced. However, $f_{\textnormal{left}}(0)=2=f_{\textnormal{left}}(1)$ and $f_{\textnormal{right}}(2)=0=f_{\textnormal{right}}(1)$. One can also check that changing the definitions of $f_{\textnormal{left}},f_{\textnormal{left}}$ to contain the restriction “strictly to the left/right” does not resolve the issue.
@@
$B=\{0,\alpha,\left\{2\alpha\right\},\ldots,\left\{(m-1)\alpha\right\}\}$
@@
Let $B=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\subseteq\mathbb{T}$ be a set of $m$ distinct points and let $\delta\in(0,1)$. Then the intervals of length $\delta$ are balanced with respect to $B$ if and only if the intervals of length $1-\delta$ are balanced. Moreover, if $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$, then for all $\varepsilon$ with $\lvert\varepsilon\rvert$ sufficiently small, the intervals of length $\delta$ are balanced if and only if the intervals of length $1-\delta+\varepsilon$ are balanced.
@@
The first statement is clear because the half open intervals of length $1-\delta$ correspond to the complements of the intervals of length $\delta$. If $\delta\neq\xi_{i}-\xi_{j}$ for all $i,j$, then it is clear that for $\lvert\varepsilon\rvert<\min_{i,j}\lvert\delta-(\xi_{i}-\xi_{j})\rvert$ we can modify the interval length $\delta$ to $\delta+\varepsilon$ without changing the occurring values of $([\xi,\xi+\delta)\cap B)$. ∎
@@ 5.  Properties of convergents and Ostrowski representations
The strategy for proving our main result (Theorem 2.1) in the next section will be to apply Lemma 4.3 to the points $\xi_{\ell}=\left\{\ell\alpha\right\}$ with $0\leq\ell\leq m-1$ and $\delta=\left\{n\alpha\right\}$. Therefore, the right endpoints of the intervals $\xi_{\ell},\xi_{\ell}+\delta)$ will be of the shape $\left\{\ell\alpha\right\}+\left\{n\alpha\right\}=\left\{(\ell+n)\alpha\right\}$, where $\ell+n>m-1$. In order to determine $f_{\textnormal{left}}(\ell),f_{\textnormal{right}}(\ell)$ we will need to figure out which point $\left\{j\alpha\right\}$ with $0\leq j\leq m-1$ lies closest to the left and which one closest to the right of $\left\{(\ell+n)\alpha\right\}$. We will do this using the Ostrowski representations mentioned in Section 2.
We now recall several properties of Ostrowski representations which that be useful. All lemmas in this section are probably well known to experts (except perhaps for Lemma 5.16, but this one is not very hard to see either). Rather unfortunately though, there does not seem to exist a suitable reference book. In order to provide proofs for the lemmas, we recall some basic properties of convergents first; see, e.g., [7] for a reference. Throughout the rest of this paper, $\alpha\in(0,1)$ is a fixed irrational, and all Ostrowski representations are with respect to $\alpha$. In fact, in view of Lemma 2.4 and Remark 2.3, we assume $\alpha<1/2$.
@@ 5.1.  Basic properties of convergents
As before, let $p_{k}/q_{k}$ denote the convergents to $\alpha\in(0,1/2)$. Then the numerators and denominators follow the recursions
@@
Then we have (see, e.g., [7, p. 9])
(5.3) $\delta_{k}=(-1)^{k}\lVert q_{k}\alpha\rVert\quad\text{for }k\geq 0,$
@@ 5.2.  Basic properties of Ostrowski representations
The next lemma describes the fact that if $k_{0}(n)$ is large, then $n\alpha$ is close to 0 modulo $1$. In particular, this shows that indeed digits with large index only have little impact on the position of $\left\{n\alpha\right\}$. Later in this paper, we will need somewhat precise bounds for the size of $\lVert n\alpha\rVert$ if $n$ only has large digits, which is why the lemma is a bit lengthy. For quick intuition, focus on the second estimate.
@@
    $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$, provided that $L\geq 1$.
Some other, more precise estimates are the following:
@@
    $\lVert n\alpha\rVert\geq\lVert q_{L+1}\alpha\rVert+\lVert q_{N+2}\rVert$ if $n<q_{N+1}$;
@@
    $\lVert n\alpha\rVert>\lVert q_{L}\alpha\rVert+\lVert q_{L+1}\alpha\rVert$ if $b_{L}(n)\geq 2$;
@@
    $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert-\lVert q_{N+2}\rVert$, provided that $L\geq 1$ and $n<q_{N+1}$;
@@
    $\lVert n\alpha\rVert<\lVert(q_{L-1}+q_{L})\alpha\rVert$ if $b_{L}(n)\leq a_{L+1}-1$ and $L\geq 1$;
@@
    $\lVert n\alpha\rVert<\lVert n^{\prime}\alpha\rVert$ if $k_{0}(n^{\prime})=k_{0}(n)=L$ and $b_{L}(n)<b_{L}(n^{\prime})$;
@@
    $\lVert n\alpha\rVert<\lVert q_{L}\alpha\rVert$ if and only if the Ostrowski representation of $n$ is of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+1+2t}>0$;
  7. (9)
    $\lVert n\alpha\rVert\geq\lVert q_{L}\alpha\rVert+\lVert q_{N}\alpha\rVert$ if $q_{L}<n<q_{N+1}$ and $n$ is not of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+1+2t}>0$.
@@
All statements follow pretty straightforwardly from (5.3), (5.7), and the rules for the digits of Ostrowski representations. See also, for example, [7, p. 24–25, Lemma 1 and Theorem 1]. ∎
The next lemma follows from similar arguments; see, for example, [3, Lemma 4.1] for a reference.
@@
Let $\alpha\in(0,1/2)$ and $n\geq 1$ and assume that $k_{0}(n)\geq 1$. Then we have
@@
Lemma 5.2 cannot be extended to $k_{0}(n)\geq 0$, so when the smallest allowed digit in the representation of $n$ shows up, then $\left\{n\alpha\right\}$ might lie on “the wrong side of 1/21/2”. Indeed, if $k_{0}(n)=0$, we might have $\left\{n\alpha\right\}>1/2$ if $b_{0}(n)$ is large, even though from the parity of $k_{0}(n)$ we would expect $\left\{n\alpha\right\}<1/2$. The next lemma will be useful when dealing with such exceptional cases.
@@
Note that the implied statement is equivalent to $\lVert(n+1)\alpha\rVert<\lVert n\alpha\rVert$.
@@
If $n=\sum_{k=0}^{N}b_{k}q_{k}$ with $b_{0}>0$, then we get from (5.3) and (5.7) that
@@
Thus, $\left\{n\alpha\right\}+\left\{\alpha\right\}<1$ and so $\left\{(n+1)\alpha\right\}=\left\{n\alpha\right\}+\left\{\alpha\right\}>\left\{n\alpha\right\}$, as desired. ∎
@@ 5.3.  One-sided approximations
The overload of notation is justified by the fact that we will go on to focus on one of $\langle\cdot\rangle_{\textnormal{left}}$, $\langle\cdot\rangle_{\textnormal{right}}$ at a time; see also the lemma below. Moreover, in Section 6 we will consider the functions $f_{\textnormal{left}},f_{\textnormal{right}},x_{\textnormal{left}},x_{\textnormal{right}}$, and the definition above will be compatible with those functions. Note that if $\{\langle\cdot\rangle_{*},\langle\cdot\rangle_{**}\}=\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$, we have
@@
$\langle\xi_{1}\rangle_{*}+\langle\xi_{2}\rangle_{*}<1\quad\iff\quad\langle\xi_{1}+\xi_{2}\rangle_{*}=\langle\xi_{1}\rangle_{*}+\langle\xi_{2}\rangle_{*};$
$\langle\xi_{1}\rangle_{*}\geq\langle\xi_{2}\rangle_{*}\quad\iff\quad\langle\xi_{1}-\xi_{2}\rangle_{*}=\langle\xi_{1}\rangle_{*}-\langle\xi_{2}\rangle_{*}.$
@@
$\langle\xi\rangle_{*}<\langle\xi+\xi_{1}\rangle_{*}<\langle\xi+\xi_{2}\rangle_{*}\quad\implies\quad\langle\xi_{1}\rangle_{*}<\langle\xi_{2}\rangle_{*}.$
We start with a simple lemma for the situation where $\langle n\alpha\rangle_{*}>1/2$.
@@
Let $\langle\cdot\rangle_{*}\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$. If $\langle n\alpha\rangle_{*}>1/2$, then either $\langle(n-1)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$ or $\langle(n+1)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$.
@@
Since either $\langle\alpha\rangle_{*}<1/2$ or $\langle-\alpha\rangle_{*}<1/2$, we have that $\langle\pm\alpha\rangle_{*}<1/2<\langle n\alpha\rangle_{*}$ for one choice of $\pm$. Then $\langle(n\mp 1)\alpha\rangle_{*}=\langle n\alpha\rangle_{*}-\langle\pm\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$. ∎
@@
(5.8) $\langle\cdot\rangle_{*}:=\begin{cases}\langle\cdot\rangle_{\textnormal{left}},&\text{if }\lVert n\alpha\rVert=\left\{n\alpha\right\};\\ \langle\cdot\rangle_{\textnormal{right}},&\text{if }\lVert n\alpha\rVert=1-\left\{n\alpha\right\}.\end{cases}$
In other words, we choose $\langle\cdot\rangle_{*}$ such that $\langle n\alpha\rangle_{*}=\lVert n\alpha\rVert<1/2$.
We rephrase Lemma 5.2 in terms of $\langle\cdot\rangle_{*}$:
@@
Let $\alpha\in(0,1/2)$ and $n,n^{\prime}\geq 1$ with $k_{0}(n),k_{0}(n^{\prime})\geq 1$. Set $\langle\cdot\rangle_{*}$ as in (5.8). Then
$\langle n^{\prime}\alpha\rangle_{*}<1/2\quad\iff\quad k_{0}(n)\equiv k_{0}(n^{\prime})\pmod{2}$.
The next two lemmas describe how we can compare $\langle n\alpha\rangle_{*}$ and $\langle n^{\prime}\alpha\rangle_{*}$ by looking at the first nonzero digits in $n,n^{\prime}$.
@@
Let $\alpha\in(0,1/2)$ and $n\geq 1$ and set $\langle\cdot\rangle_{*}$ as in (5.8). Then for all $n^{\prime}\geq 1$ we have
$k_{0}(n^{\prime})<k_{0}(n)\quad\implies\quad\langle n\alpha\rangle_{*}<\langle n^{\prime}\alpha\rangle_{*}.$
@@
If $k_{0}(n)=0$ , the implication is trivial.
Assume next that $k_{0}(n)\geq 2$. Then if $k_{0}(n^{\prime})=k_{0}(n)-1$, we have from Lemma 5.5 that $\langle n^{\prime}\alpha\rangle_{*}>1/2>\langle n\alpha\rangle_{*}$. If $k_{0}(n^{\prime})\leq k_{0}(n)-2$, then we obtain $\langle n^{\prime}\alpha\rangle_{*}>\langle n\alpha\rangle_{*}$ from Lemma 5.1(1, 2).
We are left with the case $k_{0}(n)=1$. Lemma 5.2 implies that $\langle\cdot\rangle_{*}=\langle\cdot\rangle_{\textnormal{right}}$. Assume $k_{0}(n^{\prime})=0$. If $\left\{n^{\prime}\alpha\right\}<1/2$, then $\langle n^{\prime}\alpha\rangle_{\textnormal{right}}>1/2>\langle n\alpha\rangle_{\textnormal{right}}$, as desired. If $\left\{n^{\prime}\alpha\right\}>1/2$, then as in the proof of Lemma 5.3 we see that $\left\{n^{\prime}\alpha\right\}<1-\left\{\alpha\right\}$, and so $\langle n^{\prime}\alpha\rangle_{\textnormal{right}}>\lVert\alpha\rVert=\lVert q_{0}\alpha\rVert>\lVert n\alpha\rVert=\langle n\alpha\rangle_{\textnormal{right}}$, where we used Lemma 5.1(2). ∎
@@
Let $\alpha\in(0,1/2)$ and $n,n^{\prime}\geq 1$ with $k_{0}(n)=k_{0}(n^{\prime})=L$. Assume $L\geq 1$ and set $\langle\cdot\rangle_{*}$ as in (5.8). Then
$b_{L}(n)<b_{L}(n^{\prime})\quad\implies\quad\langle n\alpha\rangle_{*}<\langle n^{\prime}\alpha\rangle_{*}.$